17. Given the following supply and demand equations
QD = 100 – 5P
QS = 10 + 5P
a) Determine the equilibrium price and quantity.
b) If the government sets a minimum price of $10 per unit, how many units would be supplied and how many would be demanded?
c) If the govt. sets a maximum price of $5 per unit, how many units would be supplied and how many would be demanded?
d) If the demand increases to
Q’D = 200 – 5P
Determine the new equilibrium price and quantity.
1
Expert's answer
2020-11-12T17:24:29-0500

a) "QD=QS" ; "100-5P=10+5P" ; "P=9;Q=55"

b) "QD=50;QS=60"

c) "QD=75;QS=35"

d) "200-5P=10+5P" ; "P=19;Q=105"